OpenStudy (anonymous):

(-7v^2-7v-7)-(-4v^2+4v+8)+(-3v^2+5v-9)

pooja195 (pooja195):

Remove the parentheses and distribute the subtraction signs to get this −7v2+−7v+−7+4v2+−4v+−8+−3v2+5v+−9 then combine like terms

OpenStudy (anonymous):

you try this one :)

OpenStudy (anonymous):

-6v^2-6v-24

OpenStudy (texaschic101):

(-7v^2-7v-7)-(-4v^2+4v+8)+(-3v^2+5v-9) -- distribute through the parenthesis -7v^2 - 7v - 7 + 4v^2 - 4v - 8 - 3v^2 + 5v - 9 now we need to gather like terms... -7v^2 + 4v^2 - 3v^2 = -10v^2 + 4v^2 = -6v^2 -7v - 4v + 5v = -11v + 5v = -6v -7 - 8 - 9 = - 24 now put it all together... -6v^2 - 6v - 24......yep....you got it :)

OpenStudy (anonymous):

nooo my medal xD